# Quantum Algorithms for the Pathwise Lasso

### v1

@@ Abstract
We present a novel quantum high-dimensional linear regression algorithm with an $\ell_{1}$-penalty based on the classical LARS (Least Angle Regression) pathwise algorithm. Similarly to available classical numerical algorithms for Lasso, our quantum algorithm provides the full regularisation path as the penalty term varies, but quadratically faster per iteration under specific conditions. A quadratic speedup on the number of features/predictors $d$ is possible by using the simple quantum minimum-finding subroutine from Dürr and Høyer (arXiv’96) in order to obtain the joining time at each iteration. We then improve upon this simple quantum algorithm and obtain a quadratic speedup both in the number of features $d$ and the number of observations $n$ by using the recent approximate quantum minimum-finding subroutine from Chen and de Wolf (ICALP’23). In order to do so, we construct, as one of our main contributions, a quantum unitary based on quantum amplitude estimation to approximately compute the joining times to be searched over by the approximate quantum minimum-finding subroutine. Since the joining times are no longer exactly computed, it is no longer clear that the resulting approximate quantum algorithm obtains a good solution. As our second main contribution, we prove, via an approximate version of the KKT conditions and a duality gap, that the LARS algorithm (and therefore our quantum algorithm) is robust to errors. This means that it still outputs a path that minimises the Lasso cost function up to a small error if the joining times are only approximately computed. Finally, in the model where the observations are generated by an underlying linear model with an unknown coefficient vector, we prove bounds on the difference between the unknown coefficient vector and the approximate Lasso solution, which generalises known results about convergence rates in classical statistical learning theory analysis.
@@ 1 Introduction
One of the most important research topics that has gotten renewed attention by statisticians and the machine learning community is high-dimensional data analysis [BEM13, BvdG11, GG08, JT09, KKMR22, WM22, ZY06]. For linear regression, this means that the number of data points is less than the number of explanatory variables (features). In machine learning parlance, this is usually referred to as overparameterisation. Solutions to such problems are, however, ill defined. Having more variables to choose from is a double-edged sword. While it gives us freedom of choice, computational cost considerations favour choosing sparse models. Some sort of regularisation in the linear model, usually in the form of a penalty term, is needed to favour sparse models.
Let us consider the regression problem defined as follows. Assume to be given a fixed design matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ and a vector of observations $\mathbf{y}\in\mathbb{R}^{n}$. The setting is such that $d\geq n$, i.e., more features than observations are given. The un-regularized $\ell_{2}$-regression problem is defined as $\min_{\bm{\beta}\in\mathbb{R}^{d}}\frac{1}{2}\|\mathbf{y}-\mathbf{X}{\bm{\beta}}\|_{2}^{2}$. Generally, different norms could be used in the penalty term. Essentially, the three popular sparse regression models reduce to three prototypes recalled below with $\ell_{0},\ell_{1},\ell_{2}$-norms and tuning parameter $\lambda$:
@@
The easiest is Ridge regression which has an analytical solution and the problem is convex, but tends to give non-sparse solutions. The problem with the $\ell_{0}$-norm is that it is not convex. The best subset selection is in general NP-hard [DK08]. Lasso regression with $\ell_{1}$-penalty is a convex relaxation of this problem and will be our central object of study throughout this paper. Lasso stands for Least Absolute Shrinkage and Selection operator and has been studied extensively in the literature because of its convexity and parsimonious solutions. Generally, the optimal tuning parameter also has to be determined from the data and as such we would like an endogenous model that determines the optimal $\bm{\beta}$ along with a specific $\lambda$. The key point is that we want a path of solutions as $\lambda$ varies [EHJT04, FHT10a, ZLZ18].
For large language models that use deep learning, overparameterisation is actually desired. In high-dimensional linear regression models, in contrast, parsimonious models are desired. The conventional statistical paradigm is to keep the number of features fixed and study the asymptotics as the sample size increases. The other paradigm is predictor selection given a large choice that fits into high-dimensional regression. Selecting which of the many variables available is part of model choice and techniques that help us in choosing the right model have significant value. Having a larger number of variables in statistical models makes it simpler to pick the right model and achieve improved predictive performance. Overparametrisation is a virtue and increases the likelihood of including the essential features in the model. Yet this should be balanced with regularisation [EHJT04, FL10, LTTT14]. Introducing an $\ell_{1}$-penalty confers a number of advantages [CDS01, HTW15, Tib96]:
@@
    The $\ell_{1}$-penalty provides models that can be interpreted in a simple way and thus have advantages compared to black boxes (deep neural networks) in terms of transparency and intelligibility [${}^{+}$19, ${}^{+}$22];
@@
    Lasso selects the true model consistently even with noisy data [ZY06], and this remains an active area of research [MOK23];
@@
Lasso-type methods are used across disciplines and in many different fields where sparsity is a desired prerequisite: genetics, compressed sensing, and portfolio optimisation [CDS01, CRT06, FNW07, PL22, UGH09, ${}^{+}$09]. Advocating Lasso-type algorithms does not preclude the fact that Lasso methods can be married with neural networks [TDK23], though it is not the scope of our work here. Lasso techniques are part of the repertoire of tools available to theoreticians and practitioners in machine learning and data analysis.
@@ 1.1 Least Angle Regression algorithm
The pathwise Lasso regression problem is the problem when the regression parameter $\lambda$ varies. The Least Angle Regression (LARS) algorithm, proposed and named by Efron et al. [EHJT04], outputs the Lasso solution $\hat{\bm{\beta}}(\lambda)$ for all $\lambda>0$. To be more precise, consider the Lasso estimator described above,
$\hat{\bm{\beta}}\in\underset{\bm{\beta}\in\mathbb{R}^{d}}{\mathrm{argmin}}\frac{1}{2}\|\mathbf{y}-\mathbf{X}{\bm{\beta}}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}\quad\text{for all}~{}\lambda>0.$ (1)
@@
$\mathcal{P}\triangleq\{\hat{\bm{\beta}}(\lambda):\lambda>0\}.$ (2)
Efron et al. [EHJT04] showed that the optimal regularisation path $\mathcal{P}$ is piecewise linear and continuous with respect to $\lambda$. This means that there exist an $m\in\mathbb{N}$ and $\infty>\lambda_{0}>\cdots>\lambda_{m-1}>\lambda_{m}=0$ and $\bm{\theta}_{0},\dots,\bm{\theta}_{m}\in\mathbb{R}^{d}$ such that [^1] Define $[n]\triangleq\{1,\dots,n\}$.
$\hat{\bm{\beta}}(\lambda)=\hat{\bm{\beta}}(\lambda_{t})+(\lambda_{t}-\lambda)\bm{\theta}_{t}\quad\text{for}~{}\lambda_{t+1}<\lambda\leq\lambda_{t}~{}~{}(t\in\{0,\dots,m-1\}).$ (3)
@@
Since the regularisation path is piecewise linear, one needs only to compute all kinks $(\lambda_{t},\hat{\bm{\beta}}(\lambda_{t}))$ for $t\in\{0,\dots,m-1\}$, from which the whole regularisation path follows by linear interpolation. That is exactly what the Least Angle Regression (LARS) algorithm proposed and named by Efron et al. [EHJT04] does (we note that a very similar idea appeared earlier in the works of Osborne et al. [OPT00a, OPT00b]). Starting at $\lambda=\infty$ (or $\lambda_{0}$) where the Lasso solution is $\hat{\bm{\beta}}(\lambda)=\mathbf{0}\in\mathbb{R}^{d}$, the LARS algorithm decreases the regularisation parameter $\lambda$ and computes the regularisation path by finding the kinks along the way with the aid of the KKT conditions. Each kink corresponds to an iteration of the algorithm. More specifically, at each iteration $t$, the LARS algorithm computes the next kink $\lambda_{t+1}$ by finding the closest point (to the previous kink) where the KKT conditions break and thus need to be updated. This is done by performing a search over the active set $\mathcal{A}\triangleq\{i\in[d]:\hat{\beta}_{i}\neq 0\}$ and another search over the inactive set $\mathcal{I}\triangleq[d]\setminus\mathcal{A}$. The search over $\mathcal{A}$ finds the next point $\lambda_{t+1}^{\rm cross}$, called crossing time, where a variable $i_{t+1}^{\rm cross}$ must leave $\mathcal{A}$ and join $\mathcal{I}$. The search over $\mathcal{I}$ finds the next point $\lambda_{t+1}^{\rm join}$, called joining time, where a variable $i_{t+1}^{\rm join}$ must leave $\mathcal{I}$ and join $\mathcal{A}$. The next kink is thus $\lambda_{t+1}=\max\{\lambda_{t+1}^{\rm cross},\lambda_{t+1}^{\rm join}\}$. The overall complexity of the LARS algorithm per iteration is $O(nd+|\mathcal{A}|^{2})$: the two searches over the active and inactive sets require $O(n|\mathcal{A}|)$ and $O(n|\mathcal{I}|)$ time, respectively, while computing the new direction $\partial\hat{\bm{\beta}}/\partial\lambda=\bm{\theta}_{t+1}$ of the regularisation path, which involves the computation of the pseudo-inverse of a submatrix of $\mathbf{X}$ specified by $\mathcal{A}$, requires $O(n|\mathcal{A}|+|\mathcal{A}|^{2})$ time. When the Lasso solution is unique, it is known [Tib13] that $|\mathcal{A}|\leq\min\{n,d\}=n$ throughout the LARS algorithm.
@@ 1.2 Our work
In the high-dimensional setting where $d\gg n$, the search over the inactive set $\mathcal{I}$, and thus the computation of the joining time, is by far the most costly step per iteration. In this work, we propose quantum algorithms for the pathwise Lasso regression problem based on the Least Angle Regression (LARS) algorithm. To the best of our knowledge, this work is the first to present a quantum version of the LARS algorithm. We propose mainly two quantum algorithms based on the LARS algorithm to speedup the computation of the joining time. Our first quantum algorithm, called simple quantum LARS algorithm, is a straightforward improvement that utilizes the well-known quantum minimum-finding subroutine from Dürr and Høyer [DH96] to perform the search over $\mathcal{I}$. We assume that the design matrix $\mathbf{X}$ is stored in a quantum-readable read-only memory (QROM) and can be accessed in time $O(\operatorname{poly}\log(nd))$. We also assume that data can be written into quantum-readable classical-writable classical memories (QRAM) of size $O(n)$, which can be later queried in time $O(\operatorname{poly}\log{n})$. The complexity of computing the joining time is now $O(n\sqrt{|\mathcal{I}|})$, where $O(n)$ is the time required to compute, via classical circuits, the joining times to be maximised over $\mathcal{I}$. The final runtime of our simple quantum algorithm is $\widetilde{O}(n\sqrt{|\mathcal{I}|}+n|\mathcal{A}|+|\mathcal{A}|^{2})$ per iteration, where we omit polylog factors in $n$ and $d$.
We then improve upon our simple quantum algorithm by approximately computing the joining times to be maximised over $\mathcal{I}$, thus reducing the factor $O(n)$ to $O(\sqrt{n})$. This is done by constructing a quantum unitary based on quantum amplitude estimation [BHMT02] that approximately computes the joining times of all the variables in the inactive set $\mathcal{I}$. We then employ the recently developed approximate quantum minimum-finding subroutine from Chen and de Wolf [CdW23] to find an estimate $\epsilon$-close to the true maximum joining time. The result is our approximate quantum LARS algorithm that achieves a quadratic speedup in both the number of features $d$ and the number of observations $n$. We also propose a sampling-based classical LARS algorithm. Our improved quantum algorithm (and the sampling-based classical algorithm), however, introduces an error in computing the joining time $\lambda_{t+1}^{\rm join}$. As one of our main contributions, we prove that the LARS algorithm is robust to errors. More specifically, we show that, by slightly adapting the LARS algorithm, it returns (and so our approximate quantum algorithm) an approximate regularisation path with error proportional to the error from computing $\lambda_{t+1}^{\rm join}$. We first define an approximate regularisation path.
@@
Given $\epsilon:\mathbb{R}_{>0}\to\mathbb{R}_{\geq 0}$, a vector $\widetilde{\bm{\beta}}\in\mathbb{R}^{d}$ is an approximate Lasso solution with error $\epsilon(\lambda)$ if
[…]
@@
Let $\epsilon\in[0,2)$. Consider an approximate LARS algorithm that returns a path $\widetilde{\mathcal{P}}=\{\widetilde{\bm{\beta}}(\lambda)\in\mathbb{R}^{d}:\lambda>0\}$ and wherein, at each iteration $t$, the joining time $\lambda_{t+1}^{\rm join}$ is approximated by $\widetilde{\lambda}_{t+1}^{\rm join}$ such that $\lambda_{t+1}^{\rm join}\leq\widetilde{\lambda}_{t+1}^{\rm join}\leq(1-\epsilon/2)^{-1}\lambda_{t+1}^{\rm join}$. Then $\widetilde{\mathcal{P}}$ is an approximate regularisation path with error $\lambda\epsilon\|\widetilde{\bm{\beta}}(\lambda)\|_{1}$.
To prove the above result, we consider an approximate version of the KKT conditions and use a duality gap for the Lasso regression. As far as we are aware, this is the first direct result on the robustness of the LARS algorithm. Compared to the work of Mairal and Yu [MY12], who also introduced an approximate LARS algorithm and employed similar techniques to analyse its correctness, the computation of the joining time in their case is exact, and errors only arise by utilising a first-order optimisation method to find an approximate solution when kinks happen to be too close.
Result 2 guarantees the correctness of all our algorithms. Their time complexities, on the other hand, are given by the theorem below and summarised in Table 1. In the following, let $\mathbf{X}^{+}$ be the Moore–Penrose inverse of $\mathbf{X}$. Also, given some matrix $\mathbf{A}$, let $\|\mathbf{A}\|_{2}$ be its the spectral norm, $\|\mathbf{A}\|_{1}=\max_{j}\sum_{i}|A_{ij}|$, and $\|\mathbf{A}\|_{\max}=\max_{i,j}|A_{ij}|$.
Result 3 (Informal version of Theorems 222627).
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ and $\mathbf{y}\in\mathbb{R}^{n}$. Assume that $\mathbf{X}$ is stored in a QROM and we have access to QRAMs and classical-samplable structures of memory size $O(n)$. Let $\delta,\epsilon\in(0,1)$, and $T\in\mathbb{N}$. Let $\alpha,\gamma\in(0,1]$ such that
[…]
@@
  • There is a quantum LARS algorithm that returns an approximate regularisation path with additive error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ and $T$ kinks with probability at least $1-\delta$ and in time
    $\widetilde{O}\left(\frac{\gamma^{-1}+\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}}{\alpha\epsilon}\sqrt{|\mathcal{I}|}+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$
@@
  • There is a classical LARS algorithm that returns an approximate regularisation path with additive error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ and $T$ kinks with probability at least $1-\delta$ and in time
    $\widetilde{O}\left(\frac{\gamma^{-2}+n\|\mathbf{X}\|_{\max}^{2}\|\mathbf{X}^{+}\|_{2}^{2}}{\alpha^{2}\epsilon^{2}}|\mathcal{I}|+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$
@@
Classical LARS 1
None $O(n|\mathcal{I}|+n|\mathcal{A}|+|\mathcal{A}|^{2})$
Simple quantum LARS 2
None $\widetilde{O}(n\sqrt{|\mathcal{I}|}+n|\mathcal{A}|+|\mathcal{A}|^{2})$
Approximate classical LARS 4
@@
$\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ $\widetilde{O}\left(\frac{\gamma^{-2}+n\|\mathbf{X}\|_{\max}^{2}\|\mathbf{X}^{+}\|_{2}^{2}}{\alpha^{2}\epsilon^{2}}|\mathcal{I}|+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$
Approximate quantum LARS 3
@@
$\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ $\widetilde{O}\left(\frac{\gamma^{-1}+\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}}{\alpha\epsilon}\sqrt{|\mathcal{I}|}+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$

### v2

@@ Abstract
We present a novel quantum high-dimensional linear regression algorithm with an $\ell_{1}$-penalty based on the classical LARS (Least Angle Regression) pathwise algorithm. Similarly to available classical numerical algorithms for Lasso, our quantum algorithm provides the full regularisation path as the penalty term varies, but quadratically faster per iteration under specific conditions. A quadratic speedup on the number of features/predictors $d$ is possible by using the simple quantum minimum-finding subroutine from Dürr and Høyer (arXiv’96) in order to obtain the joining time at each iteration. We then improve upon this simple quantum algorithm and obtain a quadratic speedup both in the number of features $d$ and the number of observations $n$ by using the approximate quantum minimum-finding subroutine from Chen and de Wolf (ICALP’23). In order to do so, we construct, as one of our main contributions, a quantum unitary based on quantum amplitude estimation to approximately compute the joining times to be searched over by the approximate quantum minimum-finding subroutine. Since the joining times are no longer exactly computed, it is no longer clear that the resulting approximate quantum algorithm obtains a good solution. As our second main contribution, we prove, via an approximate version of the KKT conditions and a duality gap, that the LARS algorithm (and therefore our quantum algorithm) is robust to errors. This means that it still outputs a path that minimises the Lasso cost function up to a small error if the joining times are only approximately computed. Furthermore, we show that, when the observations are sampled from a Gaussian distribution, our quantum algorithm’s complexity only depends polylogarithmically on $n$, exponentially better than the classical LARS algorithm, while keeping the quadratic improvement on $d$. Finally, we propose a dequantised version of our quantum algorithm that also retains the polylogarithmic dependence on $n$, albeit presenting the linear scaling on $d$ from the standard LARS algorithm.
@@ 1 Introduction
One of the most important research topics that has attracted renewed attention by statisticians and the machine learning community is high-dimensional data analysis [8, 9, 40, 53, 56, 121, 126]. For linear regression, this means that the number of data points is less than the number of explanatory variables (features). In machine learning parlance, this is usually referred to as overparameterisation. Solutions to such problems are, however, ill defined. Having more variables to choose from is a double-edged sword. While it gives us freedom of choice, computational cost considerations favour choosing sparse models. Some sort of regularisation in the linear model, usually in the form of a penalty term, is needed to favour sparse models.
Let us consider the regression problem defined as follows. Assume to be given a fixed design matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ and a vector of observations $\mathbf{y}\in\mathbb{R}^{n}$. The setting is such that $d\geq n$, i.e., more features than observations are given. Typical examples are $(n,d)=(500,12288),(7500,49152),(18737,65536)$ for image reconstructions [22, 11, 2, 93]. In portfolio optimization, an instance would be $(n,d)=(504,600),(252,486)$ [19, 84]. The un-regularized $\ell_{2}$-regression problem is defined as $\min_{\bm{\beta}\in\mathbb{R}^{d}}\frac{1}{2}\|\mathbf{y}-\mathbf{X}{\bm{\beta}}\|_{2}^{2}$. Generally, different norms could be used in the penalty term. Essentially, the three popular sparse regression models reduce to three prototypes recalled below with $\ell_{0},\ell_{1},\ell_{2}$-norms and tuning parameter $\lambda$:
@@
The easiest is Ridge regression which has an analytical solution and the problem is convex, but tends to give non-sparse solutions. The best subset selection is not convex and is NP-hard [27] in general. Lasso regression with $\ell_{1}$-penalty, on the other hand, is a convex relaxation of the best subset selection problem and will be our central object of study throughout this paper. Lasso stands for Least Absolute Shrinkage and Selection operator and has been studied extensively in the literature because of its convexity and parsimonious solutions, leading to its use across disciplines and in many different fields where sparsity is a desired prerequisite: compressed sensing [15, 20, 36], genetics [109, 118], and finance [108, 21, 81]. For large language models that use deep learning, overparameterisation is actually desired. In high-dimensional linear regression models, in contrast, parsimonious models are desired. The conventional statistical paradigm is to keep the number of features fixed and study the asymptotics as the sample size increases. The other paradigm is predictor selection given a large choice that fits into high-dimensional regression. Selecting which of the many variables available is part of model choice and techniques that help us in choosing the right model have significant value. Having a larger number of variables in statistical models makes it simpler to pick the right model and achieve improved predictive performance. Overparametrisation is a virtue and increases the likelihood of including the essential features in the model. Yet this should be balanced with regularisation [32, 35, 66]. Introducing an $\ell_{1}$-penalty confers a number of advantages [15, 51, 105]:
@@
    The $\ell_{1}$-penalty provides models that can be interpreted in a simple way and thus have advantages compared to black boxes (deep neural networks) in terms of transparency and intelligibility [77, 86];
@@
    Lasso selects the true model consistently even with noisy data [126], and this remains an active area of research [74];
@@
Generally, the optimal tuning parameter $\lambda$ has to be determined from the data and as such we would like an endogenous model that determines the optimal $\bm{\beta}$ along with a specific $\lambda$. The key point is that we require a path of solutions as $\lambda$ varies [32, 33, 125].
On the other hand, quantum computing is a new paradigm that offers faster computation for certain problems, for example, integer factorisation and discrete logarithm [97, 98]. Quantum algorithms have been proposed for unregularised linear regression [117, 115, 16, 58], Ridge regression [123, 102, 96, 23], and simple versions of Lasso regression [23]. In this paper, we propose new quantum algorithms (and also dequantised versions) for the pathwise Lasso regression based on the famous Least Angle Regression (LARS) algorithm.
@@ 1.1 Least Angle Regression algorithm
The pathwise Lasso regression problem is the problem when the regression parameter $\lambda$ varies. To be more precise, consider the Lasso estimator described above,
$\hat{\bm{\beta}}(\lambda)\in\underset{\bm{\beta}\in\mathbb{R}^{d}}{\mathrm{argmin}}\frac{1}{2}\|\mathbf{y}-\mathbf{X}{\bm{\beta}}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}\quad\text{for all}\leavevmode\nobreak\ \lambda>0.$ (1)
@@
$\mathcal{P}\triangleq\{\hat{\bm{\beta}}(\lambda):\lambda>0\}.$
Efron et al. [32] showed that the optimal regularisation path $\mathcal{P}$ is piecewise linear and continuous with respect to $\lambda$. This means that there exist an $m\in\mathbb{N}$ and $\infty>\lambda_{0}>\cdots>\lambda_{m-1}>\lambda_{m}=0$ and $\bm{\theta}_{0},\dots,\bm{\theta}_{m}\in\mathbb{R}^{d}$ such that [^1] Define $[n]\triangleq\{1,\dots,n\}$.
$\hat{\bm{\beta}}(\lambda)=\hat{\bm{\beta}}(\lambda_{t})+(\lambda_{t}-\lambda)\bm{\theta}_{t}\quad\text{for}\leavevmode\nobreak\ \lambda_{t+1}<\lambda\leq\lambda_{t}\leavevmode\nobreak\ \leavevmode\nobreak\ (t\in\{0,\dots,m-1\}).$ (2)
@@
Since the regularisation path is piecewise linear, one needs only to compute all kinks $(\lambda_{t},\hat{\bm{\beta}}(\lambda_{t}))$ for $t\in\{0,\dots,m-1\}$, from which the whole regularisation path follows by linear interpolation. That is exactly what the Least Angle Regression (LARS) algorithm proposed and named by Efron et al. [32] does (a very similar idea appeared earlier in the works of Osborne et al. [80, 79]). Starting at $\lambda=\infty$ (or $\lambda_{0}$) where the Lasso solution is $\hat{\bm{\beta}}(\lambda)=\mathbf{0}\in\mathbb{R}^{d}$, the LARS algorithm decreases the regularisation parameter $\lambda$ and computes the regularisation path by finding the kinks along the way with the aid of the KKT conditions. Each kink corresponds to an iteration of the algorithm. More specifically, at each iteration $t$, the LARS algorithm computes the next kink $\lambda_{t+1}$ by finding the closest point (to the previous kink) where the KKT conditions break and thus need to be updated. This is done by performing a search over the active set $\mathcal{A}\triangleq\{i\in[d]:\hat{\beta}_{i}\neq 0\}$ and another search over the inactive set $\mathcal{I}\triangleq[d]\setminus\mathcal{A}$. The search over $\mathcal{A}$ finds the next point $\lambda_{t+1}^{\rm cross}$, called crossing time, where a variable $i_{t+1}^{\rm cross}$ must leave $\mathcal{A}$ and join $\mathcal{I}$. The search over $\mathcal{I}$ finds the next point $\lambda_{t+1}^{\rm join}$, called joining time, where a variable $i_{t+1}^{\rm join}$ must leave $\mathcal{I}$ and join $\mathcal{A}$. The next kink is thus $\lambda_{t+1}=\max\{\lambda_{t+1}^{\rm cross},\lambda_{t+1}^{\rm join}\}$. The overall complexity of the LARS algorithm per iteration is $O(nd+|\mathcal{A}|^{2})$: the two searches over the active and inactive sets require $O(n|\mathcal{A}|)$ and $O(n|\mathcal{I}|)$ time, respectively, while computing the new direction $\partial\hat{\bm{\beta}}/\partial\lambda=\bm{\theta}_{t+1}$ of the regularisation path, which involves the computation of the pseudo-inverse of a submatrix of $\mathbf{X}$ specified by $\mathcal{A}$, requires $O(n|\mathcal{A}|+|\mathcal{A}|^{2})$ time. When the Lasso solution is unique, it is known [104] that $|\mathcal{A}|\leq\min\{n,d\}=n$ throughout the LARS algorithm.
@@ 1.2 Our work
[…]
We then improve upon our simple quantum algorithm by approximately computing the joining times to be maximised over $\mathcal{I}$, thus reducing the factor $O(n)$ to $O(\sqrt{n})$. This is done by constructing a quantum unitary based on quantum amplitude estimation [10] that approximately computes the joining times of all the variables in the inactive set $\mathcal{I}$. We then employ the approximate quantum minimum-finding subroutine from Chen and de Wolf [23] to find an estimate $\epsilon$-close to the true maximum joining time. The result is our approximate quantum LARS algorithm that achieves a quadratic speedup in both the number of features $d$ and the number of observations $n$. We also propose a sampling-based classical LARS algorithm, which can be seen as a quantisation counterpart. Our improved quantum algorithm (and the sampling-based classical algorithm), however, introduces an error in computing the joining time $\lambda_{t+1}^{\rm join}$. As one of our main contributions, we prove that the LARS algorithm is robust to errors. More specifically, we show that, by slightly adapting the LARS algorithm, it returns (and so our approximate quantum algorithm) an approximate regularisation path with error proportional to the error from computing $\lambda_{t+1}^{\rm join}$. We first define an approximate regularisation path. In the following, let $\mathbf{X}^{+}$ be the Moore–Penrose inverse of $\mathbf{X}$ and given a matrix $\mathbf{A}$, let $\|\mathbf{A}\|_{2}$ be its the spectral norm, $\|\mathbf{A}\|_{1}=\max_{j}\sum_{i}|A_{ij}|$, and $\|\mathbf{A}\|_{\max}=\max_{i,j}|A_{ij}|$.
@@
Given $\epsilon:\mathbb{R}_{>0}\to\mathbb{R}_{\geq 0}$, we say that a vector $\widetilde{\bm{\beta}}\in\mathbb{R}^{d}$ is an approximate Lasso solution with error $\epsilon(\lambda)$ if
$\frac{1}{2}\|\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}\|_{2}^{2}+\lambda\|\widetilde{\bm{\beta}}\|_{1}-\left(\min_{\bm{\beta}\in\mathbb{R}^{d}}\frac{1}{2}\|\mathbf{y}-\mathbf{X}\bm{\beta}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}\right)\leq\epsilon(\lambda)\quad\text{where}\leavevmode\nobreak\ \lambda>0.$
@@
Let $T\in\mathbb{N}$, $\epsilon\geq 0$, and $\mathbf{X}\in\mathbb{R}^{n\times d}$. For $\mathcal{A}\subseteq[d]$ of size $|\mathcal{A}|\leq T$, let $\alpha_{\mathcal{A}}>0$ be such that $\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}\leq\alpha_{\mathcal{A}}$. Consider an approximate LARS algorithm that returns a path $\widetilde{\mathcal{P}}=\{\widetilde{\bm{\beta}}(\lambda)\in\mathbb{R}^{d}:\lambda>0\}$ with at most $T$ kinks and wherein, at each iteration $t$, the joining time $\lambda_{t+1}^{\rm join}$ is approximated by $\widetilde{\lambda}_{t+1}^{\rm join}$ such that $\lambda_{t+1}^{\rm join}\leq\widetilde{\lambda}_{t+1}^{\rm join}\leq\big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\big{)}^{-1}\lambda_{t+1}^{\rm join}$. Then $\widetilde{\mathcal{P}}$ is an approximate regularisation path with error $\lambda\epsilon\|\widetilde{\bm{\beta}}(\lambda)\|_{1}$.
To prove the above result, we consider an approximate version of the KKT conditions and use a duality gap for the Lasso regression. As far as we are aware, this is the first direct result on the robustness of the LARS algorithm. Compared to the work of Mairal and Yu [78], who also introduced an approximate LARS algorithm and employed similar techniques to analyse its correctness, the computation of the joining time in their case is exact and errors only arise by utilising a first-order optimisation method to find an approximate solution when kinks happen to be too close.
Result 2 guarantees the correctness of all our algorithms. Their time complexities, on the other hand, are given by the theorem below and summarised in Table 1.
Result 3 (Informal version of Theorem 25, Theorem 29, Theorem 30).
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$, $\mathbf{y}\in\mathbb{R}^{n}$, $\delta\in(0,1)$, $\epsilon>0$, and $T\in\mathbb{N}$. Assume that $\mathbf{X}$ is stored in a QROM and we have access to QRAMs and classical-samplable structures of memory size $O(n)$. For $\mathcal{A}\subseteq[d]$ of size $|\mathcal{A}|\leq T$, assume there are $\alpha_{\mathcal{A}},\gamma_{\mathcal{A}}\in(0,1)$ such that
$\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}\leq\alpha_{\mathcal{A}}\qquad\text{and}\qquad\frac{\|\mathbf{X}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{\infty}}{\|\mathbf{X}\|_{\max}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{1}}\geq\gamma_{\mathcal{A}},$
@@
  • There is a quantum LARS algorithm that returns an approximate regularisation path with error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ and $T$ kinks with probability at least $1-\delta$ and in time
    $\widetilde{O}\left(\frac{\gamma_{\mathcal{A}}^{-1}+\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}}{(1-\alpha_{\mathcal{A}})\epsilon}\sqrt{|\mathcal{I}|}+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$
@@
  • There is a classical (dequantised) LARS algorithm that returns an approximate regularisation path with error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ and $T$ kinks with probability at least $1-\delta$ and in time
    $\widetilde{O}\left(\frac{\gamma_{\mathcal{A}}^{-2}+n\|\mathbf{X}\|_{\max}^{2}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}^{2}}{(1-\alpha_{\mathcal{A}})^{2}\epsilon^{2}}|\mathcal{I}|+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$
@@
Classical LARS 1 None $O(n|\mathcal{I}|+n|\mathcal{A}|+|\mathcal{A}|^{2})$
Simple quantum LARS 2 None $\widetilde{O}(n\sqrt{|\mathcal{I}|}+n|\mathcal{A}|+|\mathcal{A}|^{2})$
@@
Approximate classical LARS 4 $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ $\widetilde{O}\left(\frac{\gamma_{\mathcal{A}}^{-2}+n\|\mathbf{X}\|_{\max}^{2}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}^{2}}{(1-\alpha_{\mathcal{A}})^{2}\epsilon^{2}}|\mathcal{I}|+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$
@@
Approximate quantum LARS 3 $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ $\widetilde{O}\left(\frac{\gamma_{\mathcal{A}}^{-1}+\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}}{(1-\alpha_{\mathcal{A}})\epsilon}\sqrt{|\mathcal{I}|}+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$